### v1

@@ Abstract
For a graph $G=(V,E)$ with a vertex set $V$ and an edge set $E$, a function $f:V\rightarrow\{0,1,2,...,diam(G)\}$ is called a broadcast on $G$. For each vertex $u\in V$, if there exists a vertex $v$ in $G$ (possibly, $u=v$) such that $f(v)>0$ and $d(u,v)\leq f(v)$, then $f$ is called a dominating broadcast on $G$. The cost of the dominating broadcast $f$ is the quantity $\sum_{v\in V}f(v)$. The minimum cost of a dominating broadcast is the broadcast domination number of $G$, denoted by $\gamma_{b}(G)$. A multipacking is a set $S\subseteq V$ in a graph $G=(V,E)$ such that for every vertex $v\in V$ and for every integer $r\geq 1$, the ball of radius $r$ around $v$ contains at most $r$ vertices of $S$, that is, there are at most $r$ vertices in $S$ at a distance at most $r$ from $v$ in $G$. The multipacking number of $G$ is the maximum cardinality of a multipacking of $G$ and is denoted by $\operatorname{mp}(G)$. It is known that $\operatorname{mp}(G)\leq\gamma_{b}(G)$ and that $\gamma_{b}(G)\leq 2\operatorname{mp}(G)+3$ for any graph $G$, and it was shown that $\gamma_{b}(G)-\operatorname{mp}(G)$ can be arbitrarily large for connected graphs (as there exist infinitely many connected graphs $G$ where $\gamma_{b}(G)/\operatorname{mp}(G)=4/3$ with $\operatorname{mp}(G)$ arbitrarily large). For strongly chordal graphs, it is known that $\operatorname{mp}(G)=\gamma_{b}(G)$ always holds.
@@ 1 Introduction
Covering and packing problems are fundamental in graph theory and algorithms cornuejols2001combinatorial . In this paper, we study two dual covering and packing problems called broadcast domination and multipacking. The broadcast domination problem has a natural motivation in telecommunication networks: imagine a network with radio emission towers, where each tower can broadcast information at any radius $r$ for a cost of $r$. The goal is to cover the whole network by minimizing the total cost. The multipacking problem is its natural packing counterpart and generalizes various other standard packing problems. Unlike many standard packing and covering problems, these two problems involve arbitrary distances in graphs, which makes them challenging. The goal of this paper is to study the relation between these two parameters in the class of chordal graphs and $\delta$-hyperbolic graphs.
@@
For each vertex $u\in V$, if there exists a vertex $v$ in $G$ (possibly, $u=v$) such that $f(v)>0$ and $d(u,v)\leq f(v)$, then $f$ is called a dominating broadcast on $G$. The cost of the broadcast $f$ is the quantity $\sigma(f)$, which is the sum of the weights of the broadcasts over all vertices in $G$. So, $\sigma(f)=\sum_{v\in V}f(v)$. The minimum cost of a dominating broadcast in G (taken over all dominating broadcasts) is the broadcast domination number of G, denoted by $\gamma_{b}(G)$. So, $\gamma_{b}(G)=\displaystyle\min_{f\in D(G)}\sigma(f)=\displaystyle\min_{f\in D(G)}\sum_{v\in V}f(v)$, where $D(G)=$ set of all dominating broadcasts on $G$. Suppose $f$ is a dominating broadcast with $f(v)\in\{0,1\}$ $\forall v\in V(G)$, then $\{v\in V(G):f(v)=1\}$ is a dominating set on $G$. The minimum cardinality of a dominating set is the domination number which is denoted by $\gamma(G)$. An optimal broadcast or optimal dominating broadcast on a graph $G$ is a dominating broadcast with a cost equal to $\gamma_{b}(G)$. A dominating broadcast is efficient if no vertex hears a broadcast from two different vertices. So, no tower can hear a broadcast from another tower in an efficient broadcast. There is a theorem that says, for every graph there is an optimal dominating broadcast that is also efficient dunbar2006broadcasts . Define a ball of radius $r$ around $v$ by $N_{r}[v]=\{u\in V(G):d(v,u)\leq r\}$. Suppose $V(G)=\{v_{1},v_{2},v_{3},\dots,v_{n}\}$. Let $c$ and $x$ be the vectors indexed by $(i,k)$ where $v_{i}\in V(G)$ and $1\leq k\leq diam(G)$, with the entries $c_{i,k}=k$ and $x_{i,k}=1$ when $f(v_{i})=k$ and $x_{i,k}=0$ when $f(v_{i})\neq k$. Let $A=[a_{j,(i,k)}]$ be a matrix with the entries
@@
The maximum multipacking problem is the dual integer program of the above problem. Moreover, multipacking can be seen as a refinement of the classic independent set problems. A multipacking is a set $M\subseteq V$ in a graph $G=(V,E)$ such that $|N_{r}[v]\cap M|\leq r$ for each vertex $v\in V(G)$ and for every integer $r\geq 1$. The multipacking number of $G$ is the maximum cardinality of a multipacking of $G$ and it is denoted by $\operatorname{mp}(G)$. A maximum multipacking is a multipacking $M$ of a graph $G$ such that $|M|=\operatorname{mp}(G)$. If $M$ is a multipacking, we define a vector $y$ with the entries $y_{j}=1$ when $v_{j}\in M$ and $y_{j}=0$ when $v_{j}\notin M$. So,
@@
Brief Survey: Erwin erwin2004dominating ; erwin2001cost introduced broadcast domination in his doctoral thesis in 2001. Multipacking was introduced in Teshima’s Master’s Thesis teshima2012broadcasts in 2012 (also see beaudou2019multipacking ; cornuejols2001combinatorial ; dunbar2006broadcasts ; meir1975relations ). For general graphs, surprisingly, an optimal dominating broadcast can be found in polynomial-time $O(n^{6})$ heggernes2006optimal . The same problem can be solved in linear time for trees brewster2019broadcast . However, until now, there is no known polynomial-time algorithm to find a maximum multipacking of general graphs (the problem is also not known to be NP-hard). However, polynomial-time algorithms are known for trees and more generally, strongly chordal graphs brewster2019broadcast . See foucaud2021complexity for other references concerning algorithmic results on the two problems.
@@
We know that $\operatorname{mp}(G)\leq\gamma_{b}(G)$, since broadcast domination and multipacking are dual problems brewster2013new , and $\gamma_{b}(G)\leq\operatorname{rad}(G)$ erwin2001cost , where $\operatorname{rad}(G)$ is the radius of $G$. Moreover, it is known that $\gamma_{b}(G)\leq 2\operatorname{mp}(G)+3$ beaudou2019broadcast and it is a conjecture that $\gamma_{b}(G)\leq 2\operatorname{mp}(G)$ for every graph $G$ beaudou2019broadcast . If a graph $G$ has diameter 2 but no universal vertex, then $\gamma_{b}(G)=2$ (since no single vertex can dominate all other vertices using a broadcast of cost 1, but every vertex can reach every other vertex by broadcasting at distance 2) and $\operatorname{mp}(G)=1$ (since any two vertices are either adjacent or have a common neighbour). Examples of such graphs are the cycles $C_{4}$, $C_{5}$, complete bipartite graphs that are not stars, the Petersen graph, the Wagner graph, the Grötzsch graph, the Clebsch graph, etc. One can also take such a graph and blow up some of its vertices into independent sets or cliques to obtain more graphs. Thus, the conjectured upper bound of $2$ for the ratio $\frac{\gamma_{b}(G)}{\operatorname{mp}(G)}$, would be optimal for infinitely many graphs. However, no connected graph with arbitrarily large value of its multipacking number is known that reaches this ratio (in fact, no such graph is known with multipacking number 3 or more beaudou2019broadcast ; hartnell2014difference ).
Hartnell and Mynhardt hartnell2014difference constructed a family of connected graphs such that the difference $\gamma_{b}(G)-\operatorname{mp}(G)$ can be arbitrarily large and in fact, for which the ratio $\gamma_{b}(G)/\operatorname{mp}(G)=4/3$. Therefore, for general connected graphs,
$\frac{4}{3}\leq\lim_{\operatorname{mp}(G)\to\infty}\sup\Bigg{\{}\frac{\gamma_{b}(G)}{\operatorname{mp}(G)}\Bigg{\}}\leq 2.$
@@
A natural question comes to mind: What is the optimal bound on this ratio for other graph classes? It is known that $\gamma_{b}(G)=\operatorname{mp}(G)$ holds whenever $G$ is a strongly chordal graph brewster2019broadcast . Thus, a natural class to study is the class of chordal graphs.
@@
We also generalize this study as follows. Gromov gromov1987hyperbolic introduced $\delta$-hyperbolicity measures in order to study geometric group theory via Cayley graphs. Let $d$ be the shortest-path metric of a graph $G$. The graph $G$ is called a $\delta$-hyperbolic graph if for any four vertices $u,v,w,x\in V(G)$, the two larger of the three sums $d(u,v)+d(w,x)$, $d(u,w)+d(v,x)$, $d(u,x)+d(v,w)$ differ by at most $2\delta$. A graph class $\mathcal{G}$ is said to be hyperbolic if there exist a constant $\delta$ such that every graph $G\in\mathcal{G}$ is $\delta$-hyperbolic. Trees are 00-hyperbolic graphs, and chordal graphs are 1-hyperbolic brinkmann2001hyperbolicity . In general, the hyperbolicity is a measurement of deviation of the distance function of a graph from a tree metric. Many other interesting graph classes including co-comparability graphs corneil2013ldfs , asteroidal-triple free graphs corneil1997asteroidal , permutation graphs golumbic2004algorithmic , graphs with bounded chordality or treelength are hyperbolic chepoi2008diameters ; keil2017algorithm . See Figure 1 for a diagram representing the inclusion hierarchy between these classes. Moreover, hyperbolicity is a measure that captures properties of real-world graphs such as the Internet graph shavitt2004curvature or database relation graphs walter2002interactive . In this paper, we study the multipacking problem of hyperbolic graphs.
@@
Figure 1: Inclusion diagram for graph classes mentioned in this paper (and related ones). If a class $A$ has an upward path to class $B$, then $A$ is included in $B$. For the graphs in the gray classes, the broadcast domination number is equal to the multipacking number, but this is not true for the white classes.
@@
The question of the algorithmic complexity of computing the multipacking number of a graph has been repeatedly addressed by numerous authors, yet it has persisted as an unsolved challenge for the past decade. However, polynomial-time algorithms are known for trees and more generally, strongly chordal graphs brewster2019broadcast . Even for trees, the algorithm for finding a maximum multipacking is very non-trivial. Chordal graphs form a superclass of the class of strongly chordal graphs. In this paper, we provide a $(\frac{3}{2}+o(1))$-approximation algorithm to find a multipacking on chordal graphs.
@@
Hartnell and Mynhardt hartnell2014difference constructed a family of connected graphs $H_{k}$ such that $\operatorname{mp}(H_{k})=3k$ and $\gamma_{b}(H_{k})=4k$. There is no such construction for trees and more generally, strongly chordal graphs, since multipacking number and broadcast domination number are the same for these graph classes brewster2019broadcast . In this paper, we construct a family of connected chordal graphs $H_{k}$ such that $\operatorname{mp}(H_{k})=9k$ and $\gamma_{b}(H_{k})=10k$ (Figure 7).
@@
We also make a connection with the fractional versions of the two concepts dominating broadcast and multipacking, as introduced in brewster2013broadcast .
@@
A graph is $k$-chordal (or, bounded chordality graph with bound $k$) if it does not contain any induced $n$-cycle for $n>k$. The class of $3$-chordal graphs is the class of usual chordal graphs. Brinkmann, Koolen and Moulton brinkmann2001hyperbolicity proved that every chordal graph is $1$-hyperbolic. In 2011, Wu and Zhang wu2011hyperbolicity established that a $k$-chordal graph is always a $\frac{\lfloor\frac{k}{2}\rfloor}{2}$-hyperbolic graph, for every $k\geq 4$. A graph $G$ has treelength at most $d$ if it admits a tree-decomposition where the maximum distance between any two vertices in the same bag have distance at most $d$ in $G$ DG07 . Clearly, a graph of diameter $d$ has treelength at most $d$. Graphs of treelength at most $d$ have hyperbolicity at most $d$ chepoi2008diameters . Moreover, $k$-chordal graphs have treelength at most $2k$ DG07 . Additionally, it is known that the hyperbolicity is $1$ for the graph classes: asteroidal-triple free, co-comparability and permutation graphs wu2011hyperbolicity . These results about hyperbolicity and Proposition 3 directly establish the following.
@@ 3 An inequality linking Broadcast domination and Multipacking numbers of Chordal Graphs
Theorem 3 (erwin2001cost ; teshima2012broadcasts ).
@@
Theorem 4 (laskar1983powers ).
If $G$ is a connected chordal graph with radius $r$ and diameter $d$, then $22r\leq d+$ 2.
@@
From Theorem 2, $\big{\lceil}{\frac{d+1}{3}\big{\rceil}}\leq\operatorname{mp}(G)$ which implies that $d\leq 3\operatorname{mp}(G)-1$. Moreover, from Theorem 3 and Theorem 4, $\gamma_{b}(G)\leq r\leq\big{\lfloor}{\frac{d+2}{2}\big{\rfloor}}\leq\big{\lfloor}{\frac{(3\operatorname{mp}(G)-1)+2}{2}\big{\rfloor}}$ $=\big{\lfloor}{\frac{3}{2}\operatorname{mp}(G)+\frac{1}{2}\big{\rfloor}}$. Therefore, $\gamma_{b}(G)\leq\big{\lfloor}{\frac{3}{2}\operatorname{mp}(G)+\frac{1}{2}\big{\rfloor}}=\big{\lceil}{\frac{3}{2}\operatorname{mp}(G)\big{\rceil}}$. ∎

### v2

@@ Abstract
For a graph $G=(V,E)$ with a vertex set $V$ and an edge set $E$, a function $f:V\rightarrow\{0,1,2,...,diam(G)\}$ is called a broadcast on $G$. For each vertex $u\in V$, if there exists a vertex $v$ in $G$ (possibly, $u=v$) such that $f(v)>0$ and $d(u,v)\leq f(v)$, then $f$ is called a dominating broadcast on $G$. The cost of the dominating broadcast $f$ is the quantity $\sum_{v\in V}f(v)$. The minimum cost of a dominating broadcast is the broadcast domination number of $G$, denoted by $\gamma_{b}(G)$.
A multipacking is a set $S\subseteq V$ in a graph $G=(V,E)$ such that for every vertex $v\in V$ and for every integer $r\geq 1$, the ball of radius $r$ around $v$ contains at most $r$ vertices of $S$, that is, there are at most $r$ vertices in $S$ at a distance at most $r$ from $v$ in $G$. The multipacking number of $G$ is the maximum cardinality of a multipacking of $G$ and is denoted by $\operatorname{mp}(G)$.
It is known that $\operatorname{mp}(G)\leq\gamma_{b}(G)$ and that $\gamma_{b}(G)\leq 2\operatorname{mp}(G)+3$ for any graph $G$, and it was shown that $\gamma_{b}(G)-\operatorname{mp}(G)$ can be arbitrarily large for connected graphs. For strongly chordal graphs, it is known that $\operatorname{mp}(G)=\gamma_{b}(G)$ always holds.
@@ 1 Introduction
Covering and packing problems are fundamental in graph theory and algorithms [12]. In this paper, we study two dual covering and packing problems called broadcast domination and multipacking. The broadcast domination problem has a natural motivation in telecommunication networks: imagine a network with radio emission towers, where each tower can broadcast information at any radius $r$ for a cost of $r$. The goal is to cover the whole network by minimizing the total cost. The multipacking problem is its natural packing counterpart and generalizes various other standard packing problems. Unlike many standard packing and covering problems, these two problems involve arbitrary distances in graphs, which makes them challenging. The goal of this paper is to study the relation between these two parameters in the class of chordal graphs and $\delta$-hyperbolic graphs.
@@
For each vertex $u\in V$, if there exists a vertex $v$ in $G$ (possibly, $u=v$) such that $f(v)>0$ and $d(u,v)\leq f(v)$, then $f$ is called a dominating broadcast on $G$. The cost of the broadcast $f$ is the quantity $\sigma(f)$, which is the sum of the weights of the broadcasts over all vertices in $G$. So, $\sigma(f)=\sum_{v\in V}f(v)$. The minimum cost of a dominating broadcast in G (taken over all dominating broadcasts) is the broadcast domination number of G, denoted by $\gamma_{b}(G)$. So, $\gamma_{b}(G)=\displaystyle\min_{f\in D(G)}\sigma(f)=\displaystyle\min_{f\in D(G)}\sum_{v\in V}f(v)$, where $D(G)=$ set of all dominating broadcasts on $G$.
Suppose $f$ is a dominating broadcast with $f(v)\in\{0,1\}$ $\forall v\in V(G)$, then $\{v\in V(G):f(v)=1\}$ is a dominating set on $G$. The minimum cardinality of a dominating set is the domination number which is denoted by $\gamma(G)$.
An optimal broadcast or optimal dominating broadcast on a graph $G$ is a dominating broadcast with a cost equal to $\gamma_{b}(G)$. A dominating broadcast is efficient if no vertex hears a broadcast from two different vertices. So, no tower can hear a broadcast from another tower in an efficient broadcast. There is a theorem that says, for every graph there is an optimal dominating broadcast that is also efficient [15]. Define the ball of radius $r$ around $v$ by $N_{r}[v]=\{u\in V(G):d(v,u)\leq r\}$. Suppose $V(G)=\{v_{1},v_{2},v_{3},\dots,v_{n}\}$. Let $c$ and $x$ be the vectors indexed by $(i,k)$ where $v_{i}\in V(G)$ and $1\leq k\leq diam(G)$, with the entries $c_{i,k}=k$ and $x_{i,k}=1$ when $f(v_{i})=k$ and $x_{i,k}=0$ when $f(v_{i})\neq k$. Let $A=[a_{j,(i,k)}]$ be a matrix with the entries
@@
A multipacking is a set $M\subseteq V$ in a graph $G=(V,E)$ such that $|N_{r}[v]\cap M|\leq r$ for each vertex $v\in V(G)$ and for every integer $r\geq 1$. The multipacking number of $G$ is the maximum cardinality of a multipacking of $G$ and it is denoted by $\operatorname{mp}(G)$. A maximum multipacking is a multipacking $M$ of a graph $G$ such that $|M|=\operatorname{mp}(G)$. If $M$ is a multipacking, we define a vector $y$ with the entries $y_{j}=1$ when $v_{j}\in M$ and $y_{j}=0$ when $v_{j}\notin M$. So,
@@
The maximum multipacking problem is the dual integer program of the optimal dominating broadcast problem.
@@
Brief Survey: Erwin [16, 17] introduced broadcast domination in his doctoral thesis in 2001. Multipacking was introduced in Teshima’s Master’s Thesis [28] in 2012 (also see [2, 12, 15, 25]). For general graphs, surprisingly, an optimal dominating broadcast can be found in polynomial-time $O(n^{6})$ [22]. The same problem can be solved in linear time for trees [5]. However, until now, there is no known polynomial-time algorithm to find a maximum multipacking of general graphs (the problem is also not known to be NP-hard). However, polynomial-time algorithms are known for trees and more generally, strongly chordal graphs [5]. See [18] for other references concerning algorithmic results on the two problems.
@@
We know that $\operatorname{mp}(G)\leq\gamma_{b}(G)$, since broadcast domination and multipacking are dual problems [6], and $\gamma_{b}(G)\leq\operatorname{rad}(G)$ [17], where $\operatorname{rad}(G)$ is the radius of $G$. Moreover, it is known that $\gamma_{b}(G)\leq 2\operatorname{mp}(G)+3$ [1] and it is a conjecture that $\gamma_{b}(G)\leq 2\operatorname{mp}(G)$ for every graph $G$ [1]. If a graph $G$ has diameter 2 but no universal vertex, then $\gamma_{b}(G)=2$ (since no single vertex can dominate all other vertices using a broadcast of cost 1, but every vertex can reach every other vertex by broadcasting at distance 2) and $\operatorname{mp}(G)=1$ (since any two vertices are either adjacent or have a common neighbour). Examples of such graphs are the cycles $C_{4}$, $C_{5}$, complete bipartite graphs that are not stars, the Petersen graph, the Wagner graph, the Grötzsch graph, the Clebsch graph, etc. One can also take such a graph and blow up some of its vertices into independent sets or cliques to obtain more graphs. Thus, the conjectured upper bound of $2$ for the ratio $\frac{\gamma_{b}(G)}{\operatorname{mp}(G)}$, would be optimal for infinitely many graphs. However, no family of connected graphs with arbitrarily large value of their multipacking numbers is known that reach this ratio (in fact, no such graph is known with multipacking number 3 or more [1, 21]).
Hartnell and Mynhardt [21] constructed a family of connected graphs such that the difference $\gamma_{b}(G)-\operatorname{mp}(G)$ can be arbitrarily large and in fact, for which the ratio $\gamma_{b}(G)/\operatorname{mp}(G)=4/3$. Until recently, this was the best known construction. After the submission of the current paper, Rajendraprasad, Sani, Sasidharan and Sen [26] showed that hypercubes form an interesting example family in this regard. Indeed, it was already known that for the hypercube $H_{d}$ of dimension $d$, we have $\gamma_{b}(H_{d})=d-1$ [3], and the above authors [26] showed that $\operatorname{mp}(H_{d})\leq\frac{d}{2}+6\sqrt{2d}$. Therefore, for general connected graphs,
$\lim_{\operatorname{mp}(G)\to\infty}\sup\Bigg\{\frac{\gamma_{b}(G)}{\operatorname{mp}(G)}\Bigg\}=2.$
@@
A natural question comes to mind: What is the maximum value of this ratio for other classes of connected graphs? It is known that $\gamma_{b}(G)=\operatorname{mp}(G)$ holds whenever $G$ is a strongly chordal graph [5]. Thus, a natural class to study is the class of chordal graphs. Note that the construction by Hartnell and Mynhardt [21] as well as hypercubes, which provide known examples with a large ratio, are not chordal.
@@
We also generalize this study as follows. Gromov [20] introduced $\delta$-hyperbolicity measures in order to study geometric group theory via Cayley graphs. Let $d$ be the shortest-path metric of a graph $G$. The graph $G$ is called a $\delta$-hyperbolic graph if for any four vertices $u,v,w,x\in V(G)$, the two larger of the three sums $d(u,v)+d(w,x)$, $d(u,w)+d(v,x)$, $d(u,x)+d(v,w)$ differ by at most $2\delta$. A graph class $\mathcal{G}$ is said to be hyperbolic if there exists a constant $\delta$ such that every graph $G\in\mathcal{G}$ is $\delta$-hyperbolic. Trees are 0-hyperbolic graphs [8], and chordal graphs are 1-hyperbolic [7]. In general, hyperbolicity is a measurement of the deviation of the distance function of a graph from a tree metric. Many other interesting graph classes including co-comparability graphs [10], asteroidal-triple free graphs [11], permutation graphs [19], graphs with bounded chordality or treelength are hyperbolic [9, 23]. See Figure 1 for a diagram representing the inclusion hierarchy between these classes. Moreover, hyperbolicity is a measure that captures properties of real-world graphs such as the Internet graph [27] or database relation graphs [30]. In this paper, we study the multipacking problem of hyperbolic graphs.
@@
Figure 1: Inclusion diagram for graph classes mentioned in this paper (and related ones). If a class AA has an upward path to class $B$, then $A$ is included in $B$. For the graphs in the gray classes, the broadcast domination number is equal to the multipacking number, but this is not true for the white classes.
@@
The question of the algorithmic complexity of computing the multipacking number of a graph has been repeatedly addressed by numerous authors, yet it has persisted as an unsolved challenge for the past decade. However, polynomial-time algorithms are known for trees and more generally, strongly chordal graphs [5]. Even for trees, the algorithm for finding a maximum multipacking is very non-trivial. Chordal graphs form a superclass of the class of strongly chordal graphs. In this paper, we provide a $(\frac{3}{2}+o(1))$-approximation algorithm to find a multipacking on chordal graphs.
@@
Hartnell and Mynhardt [21] constructed a family of connected graphs $H_{k}$ such that $\operatorname{mp}(H_{k})=3k$ and $\gamma_{b}(H_{k})=4k$. There is no such construction for trees and more generally, strongly chordal graphs, since multipacking number and broadcast domination number are the same for these graph classes [5]. In this paper, we construct a family of connected chordal graphs $H_{k}$ such that $\operatorname{mp}(H_{k})=9k$ and $\gamma_{b}(H_{k})=10k$ (Figure 7).
@@
We also make a connection with the fractional versions of the two concepts dominating broadcast and multipacking, as introduced in [4].
@@
A graph is $k$-chordal (or, bounded chordality graph with bound $k$) if it does not contain any induced $n$-cycle for $n>k$. The class of $3$-chordal graphs is the class of usual chordal graphs. Brinkmann, Koolen and Moulton [7] proved that every chordal graph is $1$-hyperbolic. In 2011, Wu and Zhang [31] established that a $k$-chordal graph is always a $\frac{\lfloor\frac{k}{2}\rfloor}{2}$-hyperbolic graph, for every $k\geq 4$. A graph $G$ has treelength at most $d$ if it admits a tree-decomposition where the maximum distance between any two vertices in the same bag have distance at most $d$ in $G$ [14]. Clearly, a graph of diameter $d$ has treelength at most $d$. Graphs of treelength at most $d$ have hyperbolicity at most $d$ [9]. Moreover, $k$-chordal graphs have treelength at most $2k$ [14]. Additionally, it is known that the hyperbolicity is $1$ for the graph classes: asteroidal-triple free, co-comparability and permutation graphs [31]. These results about hyperbolicity and Proposition 3 directly establish the following.
@@ 3 An inequality linking Broadcast domination and Multipacking numbers of Chordal Graphs
Theorem 3 ([17, 28]).
@@
Theorem 4 ([24]).
If $G$ is a connected chordal graph with radius $r$ and diameter $d$, then $2r\leq d+2$.
@@
From Theorem 2, $\big\lceil{\frac{d+1}{3}\big\rceil}\leq\operatorname{mp}(G)$ which implies that $d\leq 3\operatorname{mp}(G)-1$. Moreover, from Theorem 3 and Theorem 4, $\gamma_{b}(G)\leq r\leq\big\lfloor{\frac{d+2}{2}\big\rfloor}\leq\big\lfloor{\frac{(3\operatorname{mp}(G)-1)+2}{2}\big\rfloor}$ $\big\lfloor{\frac{3}{2}\operatorname{mp}(G)+\frac{1}{2}\big\rfloor}$. Therefore, $\gamma_{b}(G)\leq\big\lfloor{\frac{3}{2}\operatorname{mp}(G)+\frac{1}{2}\big\rfloor}=\big\lceil{\frac{3}{2}\operatorname{mp}(G)\big\rceil}$. ∎